OpenStudy (anonymous):

help meeee /. The local grocery store receives shipments of fresh eggs daily. If each shipment of 63 is bound to have 7 broken eggs, then it can be predicted that there will be 27 broken eggs in a shipment of how many?

OpenStudy (raden):

first, find the probability to broken eggs = 7/63 = 1/9

OpenStudy (raden):

then f_h = 27 n * 1/9 = 27 solve for n

OpenStudy (anonymous):

proportion time.... 7/63 = 27/x (7 broken eggs to 63 = 27 broken eggs to x) now just cross multiply and solve for x. (7)(x) = (63)(27) can you take it from here ?